OpenStudy (anonymous):

Natasha rides her bike (at a constant speed) for 3 hours, helped by a wind of 3 miles per hour. Pedaling at the same rate, the trip back against the wind takes 9 hours. Find find the total round trip distance she traveled.

OpenStudy (asnaseer):

lets say Natasha was peddling at a velocity of 'v', and lets call the distance she travelled out initially 'd'. then using speed = distance / time, for the outward journey:\[(v+3)=d/3\]and for the journey back:\[(v-9)=d/9\]so you have 2 equations and 2 unknowns - you should be able to solve this.

OpenStudy (asnaseer):

remember you are asked for the total round trip distance - which will be 2d.

OpenStudy (anonymous):

27 miles?

OpenStudy (asnaseer):

perfect! - well done!

OpenStudy (anonymous):

Thank you!